# Tensor global symmetries and the Stueckelberg mechanism for tensor fields

### v1

@@ 6 Exotic Stueckelberg mechanism
$S_{(2,2)}=-\frac{1}{2}\int\mathrm{d}^{5}x\int_{\text{B}}\bigg{(}\mathrm{d}\Omega\star\mathrm{d}\Omega-\Omega\star\Omega\bigg{)}\,,$ (86)
@@
$\int_{\text{B}}\Omega\star\Omega=-\frac{1}{4}\big{(}\Omega_{\mu\nu}{}^{\mu\nu}\Omega_{\kappa\lambda}{}^{\kappa\lambda}+\Omega_{\mu\nu}{}^{\kappa\lambda}\Omega^{\mu\nu}{}_{\kappa\lambda}-4\Omega_{\mu\nu}{}^{\mu\kappa}\Omega^{\lambda\nu}{}_{\lambda\kappa}\big{)}\,.\qquad$ (87)
@@
$\delta\Omega_{\mu\nu|\kappa\lambda}=\partial_{[\mu}\varepsilon_{\kappa\lambda|\nu]}+\partial_{[\kappa}\varepsilon_{\mu\nu|\lambda]}+\partial_{[\mu}\partial_{[\kappa}s_{\nu]\lambda]}$ (88)
@@
$\mathring{F}=\Omega-\mathrm{d}\widetilde{t}-\widetilde{\mathrm{d}}t-\mathrm{d}\widetilde{\mathrm{d}}h$ (89)
@@
$S_{(2,2)}^{\text{St}}$ $=$ $-\frac{1}{2}\int\mathrm{d}^{5}x\int_{\text{B}}\bigg{(}\mathrm{d}\Omega\star\mathrm{d}\Omega-2\mathrm{d}\widetilde{t}\star\mathrm{d}\widetilde{t}-\Omega\star\Omega\bigg{)}$ (90)
@@
$\int_{\text{B}}2\Omega\star\widetilde{\mathrm{d}}t=$ $~{}\frac{1}{4}(\Omega_{\mu\nu}{}^{\mu\nu}\partial_{\kappa}t^{\kappa\lambda}{}_{\lambda}-2\Omega_{\mu\nu}{}^{\mu\kappa}\partial_{\kappa}t^{\nu\lambda}{}_{\lambda}+2\Omega_{\mu\nu}{}^{\mu\kappa}\partial_{\lambda}t^{\nu\lambda}{}_{\kappa}+\Omega_{\mu\nu}{}^{\kappa\lambda}\partial_{\kappa}t^{\mu\nu}{}_{\lambda})\,,$ (91)
@@
$+\Omega_{\mu\nu}{}^{\kappa\lambda}\partial_{\kappa}\partial^{\mu}h_{\lambda}{}^{\nu}+\frac{1}{2}\Omega_{\mu\nu}{}^{\mu\nu}\Box h_{\kappa}{}^{\kappa}-\frac{1}{2}\Omega_{\mu\nu}{}^{\mu\nu}\partial_{\kappa}\partial_{\lambda}h^{\kappa\lambda})\,.$ (92)
@@
$\Omega_{\mu\nu|\kappa\lambda}\to\widehat{\Omega}_{\mu\nu|\kappa\lambda}=\Omega_{\mu\nu|\kappa\lambda}+2\eta_{[\mu[\kappa}h_{\nu]\lambda]}$ (93)
The reverse approach of gauging proceeds in the same way as before, starting from the suitable Lagrangian for the massless fields read off (90) and identifying the conserved and gauge invariant $(2,2)$ current $J_{\mu\nu|\kappa\lambda}$ read off the couplings (92). This is given as
$J_{\mu\nu|\kappa\lambda}=J_{\mu\nu|\kappa\lambda}^{(t)}+J_{\mu\nu|\kappa\lambda}^{(h)}\,,$ (94)
@@
$J_{\mu\nu|\kappa\lambda}^{(t)}$ $=$ $\partial_{[\kappa}t_{\mu\nu|\lambda]}-2\eta_{[\mu[\lambda}\partial^{\rho}t_{\nu]\rho|\kappa]}+2\eta_{[\mu[\lambda}\partial_{\kappa]}t_{\nu]\rho|}{}^{\rho}+\eta_{[\kappa[\mu}\eta_{\lambda]\nu]}\partial^{\rho}t_{\rho\sigma|}{}^{\sigma}\,,$ (95)
$J_{\mu\nu|\kappa\lambda}^{(h)}$ $=$ $\partial_{[\mu}\partial_{[\kappa}h_{\nu]\lambda]}+\eta_{[\mu[\lambda}\partial_{\nu]}\partial_{\kappa]}h_{\rho}{}^{\rho}+\eta_{[\mu[\lambda}\Box h_{\nu]\kappa]}-\eta_{[\mu[\lambda}\partial_{\kappa]}\partial^{\rho}h_{\nu]\rho}-\eta_{[\mu[\lambda}\partial_{\nu]}\partial^{\rho}h_{\kappa]\rho}$ (96)
@@
$J^{(h)}_{\mu\nu|\kappa\lambda}$ $=-\frac{1}{2}S_{\mu\nu|\kappa\lambda}\,,\quad$ (97)
@@
$S_{\mu\nu|\kappa\lambda}=R_{\mu\nu|\kappa\lambda}+\frac{1}{2}(\eta_{\mu\kappa}\eta_{\nu\lambda}-\eta_{\mu\lambda}\eta_{\nu\kappa})R+\eta_{\mu\lambda}R_{\nu\kappa}-\eta_{\mu\kappa}R_{\nu\lambda}+\eta_{\nu\kappa}R_{\mu\lambda}-\eta_{\nu\lambda}R_{\mu\kappa}\,,$ (98)
whose linearization appears in (96). The current satisfies the covariant divergence condition
$\nabla^{\mu}J^{(h)}_{\mu\nu|\kappa\lambda}=0\,,$ (99)
@@
$\nabla^{\lambda}R_{\mu\nu\kappa\lambda}=\nabla_{\nu}R_{\mu\kappa}-\nabla_{\mu}R_{\nu\kappa}\quad\text{and}\quad\nabla^{\mu}R_{\mu\nu}=\frac{1}{2}\nabla_{\nu}R\,.$ (100)
@@
$\eta^{\mu\kappa}J^{(h)}_{\mu\nu|\kappa\lambda}=G_{\nu\lambda}\,$ (101)
where $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ is the Einstein tensor. In other words, the tensor given in (98) is to the Einstein tensor what the Riemann tensor is to the Ricci tensor, and it is in addition covariantly conserved. This agrees with the results of the previous sections, where the trace of the current of the secondary field is a field equation.
@@
We will work with the most interesting case of the Nambu-Golstone graviton of Section 5.1. First, we note that apart from minimal coupling of the conserved current, one may also consider nonminimal couplings [8]. To see this, we note that beside the one-derivative current (68) there exist the following electric and magnetic two-derivative currents, corresponding to the Riemann tensor and its ordinary Hodge dual in the first set of indices:
$J^{\text{e}}_{\mu\nu|\rho\sigma}$ $=$ $\partial_{[\mu}\partial_{[\rho}h_{\nu]\sigma]}\,,$ (102)
$J^{\text{m}}_{\mu\nu|\rho\sigma}$ $=$ $\ast J^{\text{e}}_{\mu\nu|\rho\sigma}=\frac{1}{2}\epsilon_{\mu\nu}{}^{\kappa\lambda}\partial_{k}\partial_{[\rho}h_{\sigma]\lambda}\,.$ (103)
@@
$\mathrm{d}^{\dagger}J^{\text{e,m}}=0=\widetilde{\mathrm{d}}^{\dagger}J^{\text{e,m}}\,.$ (104)
@@
$\mathrm{d}^{\dagger}=\widetilde{\mathrm{d}}\,\text{tr}+\text{tr}\,\widetilde{\mathrm{d}}\quad\text{and}\quad\widetilde{\mathrm{d}}^{\dagger}=\mathrm{d}\,\text{tr}+\text{tr}\,\mathrm{d}\,,$ (105)
@@
To couple these currents to background fields $T^{\text{e}}$ and $T^{\text{m}}$ respectively, which are both $(2,1)$ tensor fields in this case, we need a $(2,2)$ object. This is formed naturally as $\mathrm{d}\widetilde{T}$, which arises naturally from Eq. (60), since
$\mathrm{d}\widetilde{\mathring{F}}=\mathrm{d}\widetilde{T}-\mathrm{d}\widetilde{\mathrm{d}}h-\mathrm{d}\widetilde{\mathrm{d}}\sigma\widetilde{b}\,.$ (106)
We note that in this formalism the kinetic term for the Curtright field can also be written as the Berezin integral of $\mathrm{d}\widetilde{T}\star\mathrm{d}\widetilde{T}$, which gives the same result, up to a factor, as the one of $\mathrm{d}T\star\mathrm{d}T$. Then the nonminimal couplings of $h_{\mu\nu}$ to $T^{\text{e}}_{\mu\nu|\rho}$ and $T^{\text{m}}_{\mu\nu|\rho}$ that should be added to the action (67) besides the minimal couplings (80) are
$L_{\text{e}}$ $=$ $\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{e}}\star J^{\text{e}}=\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{e}}\star\mathrm{d}\widetilde{\mathrm{d}}h\,,$ (107)
$L_{\text{m}}$ $=$ $\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{m}}\star J^{\text{m}}=\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{m}}\star\ast\,\mathrm{d}\widetilde{\mathrm{d}}h\,.$ (108)
@@
$L_{\text{e}}$ $=$ $-\partial^{[\mu}T_{\text{e}}^{\kappa\lambda|\nu]}S_{\mu\nu|\kappa\lambda}$ (109)
$L_{\text{m}}$ $=$ $-\,\epsilon^{\nu_{1}\nu_{2}\nu_{3}\nu_{4}}\partial^{\kappa}T^{\text{m}}_{\nu_{1}\nu_{2}}{}^{\lambda}\partial_{\nu_{3}}\partial_{[\kappa}h_{\lambda]\nu_{4}}\,,$ (110)
where $S_{\mu\nu|\kappa\lambda}$ is given by the linearization of (98). Since its divergence vanishes, $L_{\text{e}}$ is a total derivative, which is not the case for $L_{\text{m}}$. Under the background gauge transformation of the magnetic background field $T^{\text{m}}$, the magnetic coupling transforms to a total derivative and drops out under spacetime integration. On the other hand, $h_{\mu\nu}$ is inert under this transformation, but transforms under the electric background gauge transformation with the shift $s_{\mu\nu}$. Then we see that under this electric transformation
$\delta_{\text{e}}L_{\text{m}}=-\,\epsilon^{\nu_{1}\nu_{2}\nu_{3}\nu_{4}}\partial^{\kappa}T^{\text{m}}_{\nu_{1}\nu_{2}}{}^{\lambda}\partial_{\nu_{3}}\partial_{[\kappa}s_{\lambda]\nu_{4}}\,.$ (111)
Having identified all these, we can now improve the discussion of Section 5.2. Starting from the action (67), we can couple minimally the conserved current (79) as in (80) with $T=T^{\text{e}}$, and add the required seagull term for the electric Curtright background field. In addition, we may add to this the new couplings $L_{\text{m}}$ to the magnetic background field. The complete action of the graviton, Kalb-Ramond field and vector field coupled to both background fields is
$S[h,b,a,T_{\text{e}},T_{\text{m}}]=S[h,b,a]+\int\mathrm{d}^{4}x\,\big{(}J_{\mu\nu|\rho}T_{\text{e}}^{\mu\nu|\rho}-\epsilon_{\kappa\lambda}{}^{\mu\nu}J_{\mu\nu|\rho\sigma}\partial^{\rho}T_{\text{m}}^{\kappa\lambda|\sigma}\big{)}+S_{\text{ct}}[T_{\text{e}}]\,,\,\,\,\,$ (112)
where $S[h,b,a]$ is the action (67) (linearized Einstein-Hilbert plus kinetic term of the Kalb-Ramond field), $J_{\mu\nu|\rho}$ is given by (79), $J_{\mu\nu|\kappa\lambda}$ is the magnetic current (103), and $S_{\text{ct}}[T_{\text{e}}]$ is the counterterm term for the electric background field, given by the third line of (54). This action is invariant under magnetic background gauge transformations,
$\delta_{\text{m}}S=0\,,$ (113)
@@
$\delta_{\text{e}}S$ $=$ $\int\mathrm{d}^{4}x\,\epsilon_{\kappa\lambda}{}^{\mu\nu}\partial^{\rho}T_{\text{m}}^{\kappa\lambda|\sigma}\partial_{\mu}\partial_{[\sigma}s_{\rho]\nu}$ (114)
@@
where in the second line we performed a partial integration to bring it to a form directly comparable to the analogous one for $p$-forms in Eq. (19). This constitutes the ’t Hooft anomaly for the graviton and it would appear in the partition function. As long as $T^{\text{m}}$ is a source and not a dynamical field, we can circumvent this by turning this background field off. The anomaly tells us though that we cannot promote both background fields to dynamical fields simultaneously.

### v2

@@ 6 Exotic Stueckelberg mechanism
$S_{(2,2)}=-\frac{1}{2}\int\mathrm{d}^{5}x\int_{\text{B}}\bigg{(}\mathrm{d}\Omega\star\mathrm{d}\Omega-\Omega\star\Omega\bigg{)}\,,$ (6.1)
@@
$\int_{\text{B}}\Omega\star\Omega=-\frac{1}{4}\big{(}\Omega_{\mu\nu}{}^{\mu\nu}\Omega_{\kappa\lambda}{}^{\kappa\lambda}+\Omega_{\mu\nu}{}^{\kappa\lambda}\Omega^{\mu\nu}{}_{\kappa\lambda}-4\Omega_{\mu\nu}{}^{\mu\kappa}\Omega^{\lambda\nu}{}_{\lambda\kappa}\big{)}\,.\qquad$ (6.2)
@@
$\delta\Omega_{\mu\nu|\kappa\lambda}=\partial_{[\mu}\varepsilon_{\kappa\lambda|\nu]}+\partial_{[\kappa}\varepsilon_{\mu\nu|\lambda]}+\partial_{[\mu}\partial_{[\kappa}s_{\nu]\lambda]}$ (6.3)
@@
$\mathring{F}=\Omega-\mathrm{d}\widetilde{t}-\widetilde{\mathrm{d}}t-\mathrm{d}\widetilde{\mathrm{d}}h$ (6.4)
@@
$S_{(2,2)}^{\text{St}}$ $=$ $-\frac{1}{2}\int\mathrm{d}^{5}x\int_{\text{B}}\bigg{(}\mathrm{d}\Omega\star\mathrm{d}\Omega-2\mathrm{d}\widetilde{t}\star\mathrm{d}\widetilde{t}-\Omega\star\Omega\bigg{)}$ (6.5)
@@
$\int_{\text{B}}2\Omega\star\widetilde{\mathrm{d}}t=$ $~{}\frac{1}{4}(\Omega_{\mu\nu}{}^{\mu\nu}\partial_{\kappa}t^{\kappa\lambda}{}_{\lambda}-2\Omega_{\mu\nu}{}^{\mu\kappa}\partial_{\kappa}t^{\nu\lambda}{}_{\lambda}+2\Omega_{\mu\nu}{}^{\mu\kappa}\partial_{\lambda}t^{\nu\lambda}{}_{\kappa}+\Omega_{\mu\nu}{}^{\kappa\lambda}\partial_{\kappa}t^{\mu\nu}{}_{\lambda})\,,$ (6.6)
@@
$+\Omega_{\mu\nu}{}^{\kappa\lambda}\partial_{\kappa}\partial^{\mu}h_{\lambda}{}^{\nu}+\frac{1}{2}\Omega_{\mu\nu}{}^{\mu\nu}\Box h_{\kappa}{}^{\kappa}-\frac{1}{2}\Omega_{\mu\nu}{}^{\mu\nu}\partial_{\kappa}\partial_{\lambda}h^{\kappa\lambda})\,.$ (6.7)
@@
$\Omega_{\mu\nu|\kappa\lambda}\to\widehat{\Omega}_{\mu\nu|\kappa\lambda}=\Omega_{\mu\nu|\kappa\lambda}+2\eta_{[\mu[\kappa}h_{\nu]\lambda]}$ (6.8)
The reverse approach of gauging proceeds in the same way as before, starting from the suitable Lagrangian for the massless fields read off (6.5) and identifying the conserved and gauge invariant $(2,2)$ current $J_{\mu\nu|\kappa\lambda}$ read off the couplings (6.7). This is given as
$J_{\mu\nu|\kappa\lambda}=J_{\mu\nu|\kappa\lambda}^{(t)}+J_{\mu\nu|\kappa\lambda}^{(h)}\,,$ (6.9)
@@
$J_{\mu\nu|\kappa\lambda}^{(t)}$ $=$ $\partial_{[\kappa}t_{\mu\nu|\lambda]}-2\eta_{[\mu[\lambda}\partial^{\rho}t_{\nu]\rho|\kappa]}+2\eta_{[\mu[\lambda}\partial_{\kappa]}t_{\nu]\rho|}{}^{\rho}+\eta_{[\kappa[\mu}\eta_{\lambda]\nu]}\partial^{\rho}t_{\rho\sigma|}{}^{\sigma}\,,$ (6.10)
$J_{\mu\nu|\kappa\lambda}^{(h)}$ $=$ $\partial_{[\mu}\partial_{[\kappa}h_{\nu]\lambda]}+\eta_{[\mu[\lambda}\partial_{\nu]}\partial_{\kappa]}h_{\rho}{}^{\rho}+\eta_{[\mu[\lambda}\Box h_{\nu]\kappa]}-\eta_{[\mu[\lambda}\partial_{\kappa]}\partial^{\rho}h_{\nu]\rho}-\eta_{[\mu[\lambda}\partial_{\nu]}\partial^{\rho}h_{\kappa]\rho}$ (6.11)
@@
$J^{(h)}_{\mu\nu|\kappa\lambda}$ $=-\frac{1}{2}S_{\mu\nu|\kappa\lambda}\,,\quad$ (6.12)
@@
$S_{\mu\nu|\kappa\lambda}=R_{\mu\nu|\kappa\lambda}+\frac{1}{2}(\eta_{\mu\kappa}\eta_{\nu\lambda}-\eta_{\mu\lambda}\eta_{\nu\kappa})R+\eta_{\mu\lambda}R_{\nu\kappa}-\eta_{\mu\kappa}R_{\nu\lambda}+\eta_{\nu\kappa}R_{\mu\lambda}-\eta_{\nu\lambda}R_{\mu\kappa}\,,$ (6.13)
whose linearization appears in (6.11). The current satisfies the covariant divergence condition
$\nabla^{\mu}J^{(h)}_{\mu\nu|\kappa\lambda}=0\,,$ (6.14)
@@
$\nabla^{\lambda}R_{\mu\nu\kappa\lambda}=\nabla_{\nu}R_{\mu\kappa}-\nabla_{\mu}R_{\nu\kappa}\quad\text{and}\quad\nabla^{\mu}R_{\mu\nu}=\frac{1}{2}\nabla_{\nu}R\,.$ (6.15)
@@
$\eta^{\mu\kappa}J^{(h)}_{\mu\nu|\kappa\lambda}=G_{\nu\lambda}\,$ (6.16)
where $G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R$ is the Einstein tensor. In other words, the tensor given in (6.13) is to the Einstein tensor what the Riemann tensor is to the Ricci tensor, and it is in addition covariantly conserved. This agrees with the results of the previous sections, where the trace of the current of the secondary field is a field equation.
@@
We will work with the most interesting case of the Nambu-Golstone graviton of Section 5.1. First, we note that apart from minimal coupling of the conserved current, one may also consider nonminimal couplings [9]. To see this, we note that beside the one-derivative current (5.16) there exist the following electric and magnetic two-derivative currents, corresponding to the Riemann tensor and its ordinary Hodge dual in the first set of indices:
$J^{\text{e}}_{\mu\nu|\rho\sigma}$ $=$ $\partial_{[\mu}\partial_{[\rho}h_{\nu]\sigma]}\,,$ (7.1)
$J^{\text{m}}_{\mu\nu|\rho\sigma}$ $=$ $\ast J^{\text{e}}_{\mu\nu|\rho\sigma}=\frac{1}{2}\epsilon_{\mu\nu}{}^{\kappa\lambda}\partial_{k}\partial_{[\rho}h_{\sigma]\lambda}\,.$ (7.2)
@@
$\mathrm{d}^{\dagger}J^{\text{e,m}}=0=\widetilde{\mathrm{d}}^{\dagger}J^{\text{e,m}}\,.$ (7.3)
@@
$\mathrm{d}^{\dagger}=\widetilde{\mathrm{d}}\,\text{tr}+\text{tr}\,\widetilde{\mathrm{d}}\quad\text{and}\quad\widetilde{\mathrm{d}}^{\dagger}=\mathrm{d}\,\text{tr}+\text{tr}\,\mathrm{d}\,,$ (7.4)
@@
To couple these currents to background fields $T^{\text{e}}$ and $T^{\text{m}}$ respectively, which are both $(2,1)$ tensor fields in this case, we need a $(2,2)$ object. This is formed naturally as $\mathrm{d}\widetilde{T}$, which arises naturally from Eq. (5.8), since
$\mathrm{d}\widetilde{\mathring{F}}=\mathrm{d}\widetilde{T}-\mathrm{d}\widetilde{\mathrm{d}}h-\mathrm{d}\widetilde{\mathrm{d}}\sigma\widetilde{b}\,.$ (7.5)
We note that in this formalism the kinetic term for the Curtright field can also be written as the Berezin integral of $\mathrm{d}\widetilde{T}\star\mathrm{d}\widetilde{T}$, which gives the same result, up to a factor, as the one of $\mathrm{d}T\star\mathrm{d}T$. Then the nonminimal couplings of $h_{\mu\nu}$ to $T^{\text{e}}_{\mu\nu|\rho}$ and $T^{\text{m}}_{\mu\nu|\rho}$ that should be added to the action (5.15) besides the minimal couplings (5.28) are
$L_{\text{e}}$ $=$ $\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{e}}\star J^{\text{e}}=\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{e}}\star\mathrm{d}\widetilde{\mathrm{d}}h\,,$ (7.6)
$L_{\text{m}}$ $=$ $\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{m}}\star J^{\text{m}}=\int_{\text{B}}\mathrm{d}\widetilde{T}^{\text{m}}\star\ast\,\mathrm{d}\widetilde{\mathrm{d}}h\,.$ (7.7)
@@
$L_{\text{e}}$ $=$ $-\partial^{[\mu}T_{\text{e}}^{\kappa\lambda|\nu]}S_{\mu\nu|\kappa\lambda}$ (7.8)
$L_{\text{m}}$ $=$ $-\,\epsilon^{\nu_{1}\nu_{2}\nu_{3}\nu_{4}}\partial^{\kappa}T^{\text{m}}_{\nu_{1}\nu_{2}}{}^{\lambda}\partial_{\nu_{3}}\partial_{[\kappa}h_{\lambda]\nu_{4}}\,,$ (7.9)
where $S_{\mu\nu|\kappa\lambda}$ is given by the linearization of (6.13). Since its divergence vanishes, $L_{\text{e}}$ is a total derivative, which is not the case for $L_{\text{m}}$. Under the background gauge transformation of the magnetic background field $T^{\text{m}}$, the magnetic coupling transforms to a total derivative and drops out under spacetime integration. On the other hand, $h_{\mu\nu}$ is inert under this transformation, but transforms under the electric background gauge transformation with the shift $s_{\mu\nu}$. Then we see that under this electric transformation
$\delta_{\text{e}}L_{\text{m}}=-\,\epsilon^{\nu_{1}\nu_{2}\nu_{3}\nu_{4}}\partial^{\kappa}T^{\text{m}}_{\nu_{1}\nu_{2}}{}^{\lambda}\partial_{\nu_{3}}\partial_{[\kappa}s_{\lambda]\nu_{4}}\,.$ (7.10)
Having identified all these, we can now improve the discussion of Section 5.2. Starting from the action (5.15), we can couple minimally the conserved current (5.27) as in (5.28) with $T=T^{\text{e}}$, and add the required seagull term for the electric Curtright background field. In addition, we may add to this the new couplings $L_{\text{m}}$ to the magnetic background field. The complete action of the graviton, Kalb-Ramond field and vector field coupled to both background fields is
$S[h,b,a,T_{\text{e}},T_{\text{m}}]=S[h,b,a]+\int\mathrm{d}^{4}x\,\big{(}J_{\mu\nu|\rho}T_{\text{e}}^{\mu\nu|\rho}-\epsilon_{\kappa\lambda}{}^{\mu\nu}J_{\mu\nu|\rho\sigma}\partial^{\rho}T_{\text{m}}^{\kappa\lambda|\sigma}\big{)}+S_{\text{ct}}[T_{\text{e}}],$ (7.11)
where $S[h,b,a]$ is the action (5.15) (linearized Einstein-Hilbert plus kinetic term of the Kalb-Ramond field), $J_{\mu\nu|\rho}$ is given by (5.27), $J_{\mu\nu|\kappa\lambda}$ is the magnetic current (7.2), and $S_{\text{ct}}[T_{\text{e}}]$ is the counterterm term for the electric background field, given by the third line of (5.2). This action is invariant under magnetic background gauge transformations,
$\delta_{\text{m}}S=0\,,$ (7.12)
@@
$\delta_{\text{e}}S$ $=$ $\int\mathrm{d}^{4}x\,\epsilon_{\kappa\lambda}{}^{\mu\nu}\partial^{\rho}T_{\text{m}}^{\kappa\lambda|\sigma}\partial_{\mu}\partial_{[\sigma}s_{\rho]\nu}$ (7.13)
@@
where in the second line we performed a partial integration to bring it to a form directly comparable to the analogous one for $p$-forms in Eq. (3.4). This constitutes the ’t Hooft anomaly for the graviton and it would appear in the partition function. As long as $T^{\text{m}}$ is a source and not a dynamical field, we can circumvent this by turning this background field off. The anomaly tells us though that we cannot promote both background fields to dynamical fields simultaneously.